OpenStudy (anonymous):

1. [9.01] When the quadratic equation y = (x − 4)(2x + 5) is written in standard form, what is the value of the coefficient "b"? (4 points) 2 3 −3 −20 2. [9.01] Identify which equation is a quadratic equation. (4 points) y = −5x − 3 y = −4x(x + 10) + 34 y + x2 = (x − 4)(x + 3) y − 2 = (x − 2)(x2 + 1) 3. [9.01] Determine whether the graph of y = 2x2 − 5x + 8 opens up or down and whether it has a maximum or minimum point. (4 points) Down and Maximum Down and Minimum Up and Minimum Up and Maximum

OpenStudy (mertsj):

Did you multiply it out and put it in the form : \[ax^2+bx+c=y\]

OpenStudy (anonymous):

No ? @Mertsj

OpenStudy (mertsj):

Well do that first. You know FOIL, I assume...or some other method for multiplying two binomials?

OpenStudy (anonymous):

I have no idea what your saying, can I have answers?

OpenStudy (mertsj):

It is against the OS Code of Conduct to just give out answers. The entire idea of the website is to help students learn not to help them get a good grade.

OpenStudy (mertsj):

(x − 4)(2x + 5) Multiply those two binomials

OpenStudy (anonymous):

Whatever

OpenStudy (mertsj):

What class are you taking?

OpenStudy (anonymous):

Algebra 1

OpenStudy (mertsj):

(a+b)(c+d)=ac+ad+bc+bd Use that

OpenStudy (anonymous):

I don't know how dude

OpenStudy (andriod09):

@ThatHippieKid to multiply two binomials, this is what you do: \[(x − 4)(2x + 5)\]\[((x\times2x+x\times5)+(-4\times2x+-4\times5))\] then you would solve it. that's how you do it.